The loss due to an earthquake in a commercial building is modelled by a random variable X with density function f(x) =0.005(20−x), for 0 < x < 20 0, elsewhere Given that the fire loss exceeds 10, what is the probability that it exceeds 18.
1
Expert's answer
2020-05-13T19:29:52-0400

P(X>x)=∫20x0.005(20-x)dx=0.0025(20-x)2, 0<x<20

P(x>18|x>10)=P(x>18)/P(x>10)=(20-18)2/(20-10)2=0.04